40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13289     Entry time: Mon Sep 4 16:30:06 2017     In reply to: 13281     Reply to this: 13290   13291
Author: gautam 
Type: Update 
Category: LSC 
Subject: Oplev loop tweaking 

Now that the DRMI locking seems to be repeatable again, I want to see if I can improve the measured MICH noise. Recall that the two dominant sources of noise were

  1. BS Oplev loop A2L - this was the main noise between 30-60Hz.
  2. DAC noise - this dominated between ~60-300Hz, since we were operating with the de-whitening filters off.

In preparation for some locking attempts today evening, I did the following:

  1. Added steeper elliptic roll-off filters for the ITMX and ITMY Oplevs. This is necessary to allow the de-whitening filters to be turned on without railing the DAC.
  2. Modified the BS Oplev loop to also have steeper high-frequency (>30Hz) roll off. The roll-off between 15-30Hz is slightly less steep as a result of this change.
  3. Measured all Oplev loop TFs - UGFs are between 4 Hz and 5 Hz, phase margin is ~30degrees. I did not do any systematic optimization of this for today.
  4. Went into the Foton filter banks for all the coil output filters, and modified the "Output" settings to be on "Input crossing", with a "Tolerance" of 10 and a "Timeout" of 3 seconds. These settings are to facilitate smooth transition between the two signal paths (without and with coil-dewhitening). The parameters chosen were arbitrary and not optimized in any systematic manner.
  5. After making the above changes, I tried engaging the de-whitening filters on ITMX, ITMY and BS with the arms locked. In the past, I was unable to do this because of a number of issues - Oplev loop shapes and Foton settings among them. But today, the switching was smooth, the single arm locks weren't disturbed when I engaged the coil de-whitening.

Hopefully, I can successfully engage a similar transition tonight with the DRMI locked. The main difference compared to this daytime test is going to be that the MICH control signal is also going to be routed to the BS.

Tasks for tonight, if all goes well:

  1. Lock DRMI.
  2. Use UGF servos to set the overall loop gains for DRMI DoFs.
  3. Reduce PRCL->MICH and SRCL->MICH coupling.
  4. Measure loop shapes of all DRMI DoFs.
  5. Make sensing matrix measurement.
  6. Engage coil-dewhitening, download data, make NB.

Unrelated to this work: the PMC was locked near the upper rail of the PZT, so I re-locked it closer to the middle of the range.

Quote:

Surprisingly, there was no evidence of REFL55 behaving weirdly tonight, and I was able to easily lock the DRMI on 1f error signals using the recipe I've been using in the last few months.

ELOG V3.1.3-